OpenStudy (anonymous):

If f(x) = 3x + 7, solve: 2[f(x^2)-5]

jhonyy9 (jhonyy9):

\[f(x ^{2})=3x ^{2}+7\]

jhonyy9 (jhonyy9):

\[2(f(x ^{2})-5)=2(3x ^{2}+7-5)=6x ^{2}+4\]

OpenStudy (anonymous):

Thanks!

jhonyy9 (jhonyy9):

\[6x ^{2}+4=2(3x ^{2}+2)=0 so 3x ^{2}= -2 so x ^{2}= -2/3 so x_1,_2=+/- isqrt(2/3)= +/- (isqrt6)/3\]

jhonyy9 (jhonyy9):

so x_1,_2 = +/- (isqrt6)/3